# The Basis of Foot-Sortable Sock Orderings

### v1

@@ 2.1. The basis of foot-sortable sock orderings
With Observation 20 it is not hard to verify that these orderings are also minimally unsortable.
3. Algorithm
We now present a deterministic recursive algorithm (Algorithm 3) for sock sorting. On input $X$, the algorithm returns a triple $(\theta,(S,R),X^{\prime})$ where $X^{\prime}\subseteq X$ and $(S,R)=\theta(X^{\prime})$. If $X$ is sortable, then $X^{\prime}=X$ and $S=R=\emptyset$, so $\theta$ is a sorting sequence that fully sorts $X$. If $X$ is unsortable, then $X^{\prime}$ is an unsortable subordering of $X$ and $R\neq\emptyset$. Algorithm 3 closely follows the proof Theorem 1, so the proof of Theorem 1 also serves as a proof that Algorithm 3 is correct.
Algorithm 2 (what we call the greedy algorithm) is mentioned extensively in our proof. We often write $(S,R)=\phi_{G}(X)$ to mean that $(S,R)$ is the state when the greedy algorithm terminates. The greedy algorithm simply sorts the sortable color that terminates first in $SR$ until no color is sortable. In particular, if a color only appears on the top of the stack, then that color is sorted next.
At a high-level, the main idea of our sorting algorithm is as follows. Let $(S,R)=\phi_{G}(X)$ be the state after the greedy algorithm terminates. If we assume $X$ is sortable and the greedy algorithm fails, then we can always find a color $d$ such that any sorting sequence that sorts $X-d$ also sorts $X$. Therefore, we can recursively call our sorting routine on $X-d$ and use the resulting sequence to sort $X$. In most cases, this color $d$ is the last color sorted by the greedy algorithm. However, in some cases, we must backtrack in order to find such a color $d$.
For brevity, we do explicitly describe the subroutine Move which is referenced below. For a given state $(S,R)$ with sortable color $a$, the call $\textrm{Move}((S,R),a)$ returns the state $(S^{\prime},R^{\prime})$ after sorting the color $a$.
Algorithm 1 Sort using a prescribed sorting order, except we always immediately sort colors that only appear on the top of the stack. The routine returns the actual sequence of colors sorted.
1:function SortColors($X,\theta$)
2:     $(S,R)\leftarrow(\emptyset,X)$
3:     $\theta^{\prime}\leftarrow\emptyset$triangleright $\theta^{\prime}$ records the actual sorting sequence.
4:     while $R\neq\emptyset$ and $S\neq\emptyset$ do
5:         if $S\neq\emptyset$ and the top color $t$ on $S$ is not in $R$ then
6:              Append $t$ to $\theta^{\prime}$
7:              $(S,R)\leftarrow\textrm{Move}((S,R),t)$
8:         else if $\theta=\emptyset$ then return $\theta^{\prime},(S,R)$
9:         else
10:              $d\leftarrow$ first color in $\theta$
11:              Remove $d$ from $\theta$
12:              if $d$ is in $SR$ then
13:                  if $d$ is not sortable in $(S,R)$ then return $\theta^{\prime},(S,R)$
14:                  end if
15:                  Append $d$ to $\theta^{\prime}$
16:                  $(S,R)\leftarrow\textrm{Move}((S,R),d)$
17:              end if
18:         end if
19:     end while
20:     return $\theta^{\prime},(S,R)$
21:end function
Algorithm 2 When $C=\emptyset$ this is our greedy sorting algorithm. Otherwise, we assume $C\subseteq X$ and follow the greedy sorting algorithm, except we stop when a color in $C$ would be sorted next. We use the second parameter $C$ to backtrack.
1:function Greedy($X$, C)
2:     $(S,R)\leftarrow(\emptyset,X)$
3:     while $\exists d$ a sortable color in $(S,R)$ that terminates first in $SR$ do
4:         if $d\in C$ then return $\theta,(S,R)$
5:         end if
6:         $(S,R)\leftarrow$ $\textrm{Move}((S,R),d)$
7:         Append $d$ to $\theta$
8:     end while
9:     return $\theta,(S,R)$
10:end function
Algorithm 3 The sorting algorithm
1:function FullSort($X$)
2:     $\theta,(S,R)\leftarrow\textrm{Greedy}(X,\emptyset)$
3:     if $R=\emptyset$ or $S=\emptyset$ or $\exists x\in S$ s.t. $(x,R)$ is terminal then return $\theta,(S,R),X$
4:     end if
5:     Let $b$ and $t$ be the top two colors on $S$ and let $d$ be the last color in $\theta$
6:     if $(bt,S)$ is terminal then return $\textrm{SortHelper}(X,d,b)$
7:     end if
8:     $a\leftarrow$ the lowest color in $S$ s.t. $(abt,R)$ is terminal
9:     if $\exists z\in R$ s.t. either $z=b$ or $z$ is between $a$ and $b$ on $S$ then
10:         if $R$ contains $azt$ or $atz$ then return $\textrm{SortHelper}(X,d,a)$
11:         else return $\textrm{SortHelper}(X,d,z)$
12:         end if
13:     end if
14:     Let $A$ and $C$ be the strings that precede and follow the last $b$ in $X$ (so $X=AbC$ and $b\notin C$)
15:     $\theta,(S,R)\leftarrow\textrm{Greedy}(X,C)$triangleright Backtrack to before the final $b$ is pushed onto the stack
16:     $c\leftarrow$ first terminating sortable color in $(S,R)$triangleright $c$ would be sorted next greedily
17:     $d\ \leftarrow$ last color in $\theta$triangleright $d$ is the last color sorted
18:     if $ca\subseteq A$ then
19:         return $\textrm{Greedy}(X,\emptyset),X$triangleright $X$ is unsortable - return greedy output
20:     end if
21:     Let $y$ be the color immediately beneath the top color on the stack $S$
22:     return SortHelper(X, d, y)
23:end function
Algorithm 4 This is the helper routine for the sorting algorithm. When it is called, $X$, $a$, $x$, and $d$ must satisfy the hypothesis of Lemma 25. Furthermore, we also require that if Lemma 25(I) holds, then $X$ satisfies the hypothesis of either Lemma 27 or Lemma 28, so $X$ is unsortable. Note that when both Lemma 25(I) and (II) do not hold, if a sequence that fully sorts $X-d$ does not fully sort $X$, Lemma 25(IV) must hold and Lemma 25(IV) implies that $X$ is unsortable.
1:function SortHelper($X,d,a$)
2:     if $a$ is blocked by a $d$-sandwich in $X$ then return $\textrm{Greedy}(X,\emptyset),X$triangleright $X$ is unsortable
3:     end if
4:     $\theta,(S,R),X^{\prime}\leftarrow\textrm{FullSort}(X-d)$
5:     if  $R\neq\emptyset$  then return $\theta,(S,R),X^{\prime}$triangleright $X^{\prime}\subseteq X-d$ is unsortable
6:     else
7:         return $\textrm{SortColors}(X,\theta),X$triangleright $\theta$ fully sorts $X-d$, so $\theta$ fully sorts $X\iff X$ is sortable
8:     end if
9:end function
4. Prelimanary Observations and Lemmas
@@
Let $X$ be a sock ordering that, for some $n\geq 0$, contains distinct colors $t_{0},\dotsc,t_{n}$, $b$, and $c$ where the final $b$ precedes the final $c$ in $X$. Define $A$ and $C$ to be the strings such that $X=AbC$ where $b\notin C$ and $c\in C$. Let $\theta$ be a sorting sequence for $X$ that sorts a color in $C$ and let $c\in C$ be the color in $C$ that is sorted first by $\theta$. Taking $(S,R):=\theta(X)$, suppose $b\in S$ and that once $c$ is sorted, $b$ never appears as the top element on the stack. Let $t_{0},\dots,t_{n}$ (in this order) be the $n+1$ distinct colors which appear immediately above the color $b$ in the stack after $c$ is sorted. Then
@@
We will prove the claim by induction on $n$.
For the base case, assume $n=0$, so $t_{0}=t_{n}$ is the unique color that appears directly above $b$ after $c$ is sorted. Since $c$ follows the last $b$ and $b$ can never appear only as the top color on the stack after $c$ is sorted, the color $t_{0}=t_{n}$ must appear above $b$ after $c$ is sorted. Therefore, we have
@@
Now, suppose $n\geq 1$. Let $d$ be the color that is sorted that placed the color $t_{n}$ directly above the color $b$ on the stack. And let $\theta_{d}$ be the initial sequence of $\theta$ up to but not including $d$. Let $(S^{\prime},R^{\prime}):=\theta_{d}(X)$. Note, during the execution of $\theta_{d}$, the color $b$ is never the top element of the stack and that $t_{0},\dotsc,t_{n-1}$ is the sequence of colors that appear immediately above the color $b$ after $c$ is sorted. Therefore, by induction, we have
@@
Observation 23.
Let $X$ be a sock ordering that contains a sandwich. If $aba$ is the first sandwich to appear in $X$, then any color that is not $a$ or $b$ is either sortable in $X$ or is blocked by the sandwich $aba$.
@@
Let $S$ and $R$ be strings where $S$ contains at least two distinct colors and does not contain a sandwich. Let $a\in S$ where $a$ is not the top color in $S$ and let $S^{\prime}$ be formed by removing the color $a$ from $S$. If $x$ is unsortable in $(S,R)$ and $x$ is sortable in $(S^{\prime},R)$, then
@@
  • there exists a color $y$ distinct from the colors $a$ and $x$ such that $yx\subseteq R$ and $y$, $a$, and $x$ are the top three colors on the stack $S$, respectively
@@
Let $yzy$ be the first sandwich in $SR$ that blocks $x$. Since $S$ does not contain $yzy$, we have $yx\subseteq R$. Because $x$ is sortable in $S^{\prime}R$, the color $a$ is either $y$ or $z$. If $a$ is $y$, then $ax\subseteq R$, so assume $a$ is $z$. Since $yzy=yay$ does not block $x$ in $S^{\prime}R$, we must have $ya\subseteq S$. Recall that, by assumption, $a$ is not the top color in $S$, so let $w$ be the color that appears directly above $a$ in $S$. If $w\neq x$, then $ywy$ blocks $x$ in $S^{\prime}R$, so we have $w=x$. Since $x$ is sortable in $(S^{\prime},R)$ and $x$ is on $S^{\prime}$, the color $x$ must be the top color on the stack $S^{\prime}$. Therefore, $x$ is the top color on $S$.
We now have that the colors $a$ and $x$ are the top two colors in $S$, respectively, and $yx\subseteq R$. Therefore, we only need to show that the color $y$ is directly below the color $a$ in $S$. This follows from the observation that if there exists $u$ that is distinct from $y$, $a$, or $x$ that is between $y$ and $a$ in $S$, then the sandwich $yuy$ blocks $x$ in $S^{\prime}R$, a contradiction. ∎
The following lemma establishes some general conditions that imply an unsortable sock ordering must either contain a proper subordering that is unsortable or contain $I_{1,n}$ for some $n\geq 0$. This lemma is a crucial piece of our recursive algorithm, because if (I) and (IV) do not hold, then we can determine if a sock ordering is sortable and how to sort it by considering a proper subordering.
@@
Let $X$ be a sock ordering and let $a$, $x$, and $d$ be distinct colors in $X$. Let $\overline{\theta}$ be a nonempty sorting sequence for $X$ such that $\overline{\theta}(X)=(S,R)$, where $ax\subseteq S$, $a\in R$, and $d$ is the last color sorted in $\overline{\theta}$. Further suppose that there exists $S^{\prime}\subseteq S$ and $R^{\prime}\subseteq R$ such that $ax\subseteq S^{\prime}$ and $a$ is the only good sortable color in $S^{\prime}R^{\prime}$. One of the following is true:
@@
    $X-d$ is sortable and if a sequence $\theta$ fully sorts $X-d$, then $\theta$ also fully sorts $X$; or
@@
Assume that (I), (II), and (III) do not hold. We will show that (IV) must hold. Let $X=YDW$, where the first element of $D$ is the first $d$ in $X$ and the final element of $D$ is the final $d$ in $X$. Note that $R\subseteq W$, so $a\in W$. Because there are no $d$-sandwiches that block $a$ in $X$, the only colors that could appear in $D$ are $a$ and $d$. Since $d$ is sorted by $\overline{\theta}$ but $a$ and $x$ are not, there is no sandwich that consists only of $a$ and $x$ that blocks $d$ in $X$. Therefore, $D$ cannot contain $a$ as otherwise a sandwich of the form $axa$ would block $d$ in $X$. As such, $d$ is the only color that appears in $D$. Since $d$ is the last color in $\overline{\theta}$, this implies that $Y\supseteq S\supseteq S^{\prime}\supseteq ax$.
@@
Since, by assumption, the statements (II) and (III) are both false, there exists a sorting sequence $\theta$ that fully sorts $X-d=YW$ such that $\theta$ does not fully sort $X=YDW$. Let $\theta_{a}$ be the list of initial colors in $\theta$ up to, but not including, $a$ (so $\theta_{a}$ is empty if and only if $\theta$ begins with $a$). Note that no color in $S^{\prime}R^{\prime}\subseteq X$ could appear before $a$ in $\theta$ because $a$ is the only good sortable color in $S^{\prime}R^{\prime}$. This means that no color that appears after $ax$ in $X$ is in $\theta_{a}$. To see this, note that sorting such a color would put $ax$ onto the stack, but that would make it impossible to sort $a$ before sorting $x$. In particular, this means that no color that appears in $W$ is in $\theta_{a}$. Therefore, $a$ is the color in $W$ that is sorted first by $\theta$ and $\theta_{a}$ is a sorting sequence for $X=YDW$. Furthermore, if we let $(S_{1},R^{\prime}_{1})=\theta_{a}(YW)$, then $(S_{1},R_{1})=\theta_{a}(YDW)$ where $R^{\prime}_{1}$ is formed by removing $d$ from $R_{1}$. Therefore, because $a$ is sortable in $(S_{1},R^{\prime}_{1})$ and no $d$-sandwich blocks $a$ in $X$, the color $a$ is also sortable in $(S_{1},R_{1})$. Recall that no color that appears after $ax$ is in $\theta_{a}$. This, with the fact that $axDa\subseteq SDR\subseteq X$, implies that $xDa\subseteq R_{1}$. Therefore, sorting $a$ in $(S_{1},R_{1})$ pushes $xD$ onto the stack.
@@
Let $z$ be the first color in $\theta$ that causes a sandwich to appear in the stack when $\theta$ is applied to $X$. Such a color $z$ must exist because $\theta$ does not fully sort $X$. Let $\theta_{z}$ be the list of initial colors in $\theta$ up to but not including $z$. Note that by our previous arguments, $a$ must appear in $\theta_{z}$ and $a$ is the color in $W$ that is sorted first by $\theta_{z}$. Furthermore, since $xD$ is pushed onto the stack when $a$ is sorted and $d\notin\theta_{z}$, if we let $(S^{\prime}_{2},R_{2})=\theta_{z}(YW)$ we have $(S_{2},R_{2})=\theta_{z}(YDW)$ where $xD\subseteq S_{2}$ and $S^{\prime}_{2}=S_{2}-d$. Note that while applying $\theta_{z}$ to $X$, after $a$ is sorted the color $d$ can never appear as the top element of the stack since $\theta$ fully sorts $X-d$ but does not not fully sort $X$. Let $t_{0},\dotsc,t_{n}$ be the $n+1$ colors that appear immediately above $d$ on the stack after $a$ is sorted by $\theta_{z}$. Since $X=YDW$ and $D$ consists solely of one or more copies of $d$, Lemma 21 implies that $dI(a,t_{0},\dotsc,t_{n})R_{2}\subseteq dW$. Recall that $xD\subseteq S_{2}$. Therefore, there exists a color $y$ that appears directly below $D$ in $S_{2}$. (We could have $y=x$.) Since $axDa\subseteq X$, we have $ayDa\subseteq X$, so $aydI(a,t_{0},\dotsc,t_{n})R_{2}\subseteq YdW\subseteq X$.

### v2

@@ 2.2. The basis of foot-sortable sock orderings
With Observation 2 it is not hard to verify that these orderings are also minimally unsortable.
As we have now specified a collection $\Gamma$ of minimally unsortable sock orderings, to prove that $\Gamma$ is the basis of foot-sortable sock orderings, we only need the following theorem.
Theorem 3.
If $X$ is an unsortable sock ordering, then a subpattern of $X$ is in $\Gamma$.
2.3. Algorithm
Our proof of Theorem 3 is inherently algorithmic, and we feel it is instructive to explicitly describe a deterministic algorithm that, given an arbitrary sock order, either sorts the ordering using only the operations stack and unstack, or indicates that the given sock ordering is unsortable. Our proof can be used to extend this algorithm so that, when the given sock ordering is unsortable, it also identifies a subpattern which is an element of the basis, thereby justifying the indication that the sock ordering is unsortable. For simplicity, we do not explicitly describe this extension. We made no attempt to optimize the running time of this algorithm. (Note that, as referenced in the introduction, independent of our work, Yu has already described a fast algorithm [6].)
The algorithm first attempts to sort using the following simple greedy algorithm. This greedy algorithm is the only part of this section that is referenced in the other sections of this paper.  Greedy algorithm - Input: A sock ordering $X$.
  1. (G.1)
    While the current state is not terminal, do the following.
    1. (G.1.1)
      Sort the sortable color which terminates first. That is, if there is a color that only appears on the top of the stack, then sort that color, and otherwise sort the sortable color on the right which terminates first.
The main algorithm also makes use of the following helper routine. The input to the helper routine is a sock ordering $X$ and colors $d$ and $a$ where we assume the input satisfies the hypothesis of Lemma 8 in the following sense: There is a way to sort $X$ to a state $(S,R)$ so that
  • $d$ was the color sorted immediately before entering the state $(S,R)$;
  • for some substack $S^{\prime}\subseteq S$ and subordering $R^{\prime}\subseteq R$, the color $a$ is beneath the top color on $S^{\prime}$ and the color $a$ is the only good sortable color in $S^{\prime}R^{\prime}$. (In particular, this implies that $(S^{\prime},R^{\prime})$, and hence $(S,R)$, is unsortable.)
When this routine is called, the colors $a$ and $d$ will also have the property that if $a$ is blocked by a $d$-sandwich in $X$ (i.e. Lemma 8(I) holds), then (by either Lemma 10 or Lemma 11) a subpattern of $X$ is in $\Gamma$, so $X$ is unsortable.
 Helper routine - Input: A sock ordering $X$ and colors $d$ and $a$.
  1. (H.1)
    If $a$ is blocked by a $d$-sandwich, then $X$ is unsortable.
  2. (H.2)
    Otherwise, we recursively run the main algorithm on the smaller ordering $X-d$.
  3. (H.3)
    If it fails, then by our recursive assumption $X-d$ is unsortable, so $X$ is unsortable.
  4. (H.4)
    Otherwise, $X-d$ is sortable and we have the following two possibilities.
    1. (H.4.1)
      When we attempt to sort $X$ using the same sequence of colors that fully sorted $X-d$, there is a stage in which the color $d$ only appears as the top color of the stack, which means that we can fully sort $X$ by sorting $d$ at this stage and then continuing with the sequence that was used to sort $X-d$.
    2. (H.4.2)
      $X$ is unsortable because $I_{1,n}\in\Gamma$ is a subpattern of $X$.
We now present the main algorithm. It closely follows the proof Theorem 3 in Section 2.5, so the proof of Theorem 3 also serves as a proof that this algorithm is correct.
 Main Algorithm - Input: A sock ordering $X$.
  1. (M.1)
    Let $(S,R)$ be the output of the greedy algorithm performed on $X$.
  2. (M.2)
    If the greedy algorithm succeeds (i.e. $S=R=\emptyset$), then we are done.
  3. (M.3)
    If $S=\emptyset$ or there is a color $x\in S$ on the stack such that $(x,R)$ is terminal, then $X$ has a trivially unsortable subpattern, so $X$ is unsortable.
  4. (M.4)
    Otherwise, there are at least two colors on the stack, so let $t$ be the top color and let $b$ be the color immediately beneath it on the stack. We can also let $d$ be the final color sorted by the greedy algorithm.
  5. (M.5)
    If $(bt,R)$ is terminal, then we can finish by running the helper routine with input $X$, $d$, and $b$ (c.f. Lemma 11 and Case 2 in Theorem 3).
  6. (M.6)
    Otherwise, let $a$ be the lowest color on the stack such that $(abt,R)$ is terminal. (Such a color exists by Lemma 12.) We must have $at\subseteq R$. (This follows from Claim 14 in the proof of Theorem 3.)
  7. (M.7)
    If there exists a color $z$ that is both in $R$ and above $a$ and below or equal to $b$ on the stack $S$, then do the following (c.f. Lemma 10 and Case 3.1 in Theorem 3).
    1. (M.7.1)
      If $azt\subseteq R$ or $atz\subseteq R$, then we can finish by running the helper routine with input $X$, $d$, and $a$.
    2. (M.7.2)
      Otherwise, $zat\subseteq R$, and we can finish by running the helper routine with input $X$, $d$, and $z$.
  8. (M.8)
    Otherwise, we have $b\notin R$, so if we backtrack to the state $(S^{\prime},R^{\prime})$ immediately before the final $b$ was pushed onto the stack the following holds:
    • There exists a string $C$ with $b\notin C$ and strings $A$ and $A^{\prime}$ such that $X=AbC$ and $R^{\prime}=A^{\prime}bC$ and $S^{\prime}A^{\prime}\subseteq A$.
    • The next color sorted by the greedy algorithm (i.e. the color that pushes the final $b$ onto the stack) is a color $c\in C$.
    • For some nonnegative integer $n$, there are $n+1$ distinct colors $t_{0},\dotsc,t_{n}$ that are distinct from $b$, $c$ and $a$, such that $t_{0}cI(t_{0},\dotsc,t_{n})at_{n}\subseteq C$. (This follows from Lemma 4.)
  9. (M.9)
    Therefore, if $ca\subseteq A$, then $I_{1,n}\in\Gamma$ is a subpattern of $X$, so $X$ is unsortable.
  10. (M.10)
    Otherwise, there are at least three distinct colors on the stack $S^{\prime}$. (This follows from Claim 15 in the proof of Theorem 3.)
  11. (M.11)
    In particular, this implies that there is a color $d^{\prime}$ that was sorted immediately before the greedy algorithm reached the state $(S^{\prime},R^{\prime})$.
  12. (M.12)
    Furthermore, if we let $y$ be the color immediately beneath the top color on the stack $S^{\prime}$, we can finish by running the helper routine with input $X$, $d^{\prime}$, and $y$ (c.f. Lemma 10 and Claim 15 in the proof of Theorem 3).
2.4. Observations and Lemmas
@@
Let $X$ be a sock ordering that, for some $n\geq 0$, has distinct colors $b$, and $c$ where the final $b$ precedes the final $c$ in $X$. Define $A$ and $C$ to be the strings such that $X=AbC$ where $b\notin C$ and $c\in C$. Let $\theta$ be a nonempty sorting sequence for $X$ that sorts a color in $C$ and suppose $c$ is the color in $C$ that is sorted first by $\theta$. Taking $(S,R):=\theta(X)$, suppose $b\in S$ and that once $c$ is sorted, $b$ never appears as the top color on the stack. Suppose $t_{0},\dots,t_{n}$ (in this order) are the $n+1$ colors which appear immediately above the color $b$ in the stack after $c$ is sorted. Then
@@
Furthermore, the colors $t_{0},\dotsc,t_{n}$ are distinct and disjoint from $b$ and $c$.
@@
We will prove the first statement by induction on $n$.
For the base case, assume $n=0$, so $t_{0}=t_{n}$ is the only color that appears directly above $b$ after $c$ is sorted. Since $c$ follows the last $b$ and $b$ can never appear as the top color on the stack after $c$ is sorted, the color $t_{0}=t_{n}$ must appear above $b$ after $c$ is sorted. Therefore, we have
@@
Now, suppose $n\geq 1$. Let $d$ be the color that is sorted that placed the color $t_{n}$ directly above the color $b$ on the stack. And let $\theta_{d}$ be the initial sequence of $\theta$ up to but not including $d$. Let $(S^{\prime},R^{\prime}):=\theta_{d}(X)$. Note, during the execution of $\theta_{d}$, the color $b$ is never the top color on the stack and that $t_{0},\dotsc,t_{n-1}$ is the sequence of colors that appear immediately above the color $b$ after $c$ is sorted. Therefore, by induction, we have
@@
For the final statement, note that, for every $i\in\{0,\dotsc,n\}$, the color $t_{i}\neq b$ by definition and $t_{i}\neq c$ since it appears in the stack after $c$ is sorted. Furthermore, the colors $t_{0},\dotsc,t_{n}$ must be $n+1$ distinct colors because the only way for a color to removed from the stack is for it to be sorted.
@@
There is a natural ordering of the sandwiches that appear in a given sock ordering. If $aba$ and $xyx$ are suborderings of a sock ordering $X$ we say that $aba$ precedes $xyx$ if the final $a$ in $aba$ precedes the final $x$ in $xyx$; or when $a=x$ and the final socks in $aba$ and $xyx$ are the same, the $b$ in $aba$ precedes the $y$ in $xyx$; or when $a=x$ and $b=y$ and the last two sock in $aba$ and $xyx$ are the same, the first $a$ in $aba$ precedes the first in $xyx$.
Observation 6.
If $aba$ is the first sandwich to appear in a sock ordering $X$, then any color that is not $a$ or $b$ is either sortable in $X$ or is blocked by the sandwich $aba$.
@@
Let $S$ and $R$ be strings where $S$ is sorted and has at least two distinct colors. Let $a\in S$ where $a$ is not the top color in $S$ and let $S^{\prime}$ be formed by removing the color $a$ from $S$. If $x$ is unsortable in $(S,R)$ and $x$ is sortable in $(S^{\prime},R)$, then
@@
  • there exists a color $y$ distinct from the colors $a$ and $x$ such that $yx\subseteq R$ and $y$ is the top color, $a$ is immediately below the color $y$, and $x$ is immediately below the color $a$ on the stack $S$.
@@
Let $yzy$ be the first sandwich in $SR$ that blocks $x$. Since $S$ is sorted, $yzy$ is not a subordering of $S$, so we have $yx\subseteq R$. Because $x$ is sortable in $S^{\prime}R$, the color $a$ is either $y$ or $z$. If $a$ is $y$, then $ax\subseteq R$, so assume $a$ is $z$. Since $yzy=yay$ does not block $x$ in $S^{\prime}R$, we must have $ya\subseteq S$. Recall that, by assumption, $a$ is not the top color in $S$, so let $w$ be the color that appears directly above $a$ in $S$. If $w\neq x$, then $ywy$ blocks $x$ in $S^{\prime}R$, so we have $w=x$. Since $x$ is sortable in $(S^{\prime},R)$ and $x$ is on $S^{\prime}$, the color $x$ must be the top color on the stack $S^{\prime}$. Therefore, $x$ is the top color on $S$.
We now have that the colors $a$ and $x$ are the top two colors on $S$, respectively, and $yx\subseteq R$. Therefore, we only need to show that the color $y$ is directly below the color $a$ in $S$. This follows from the observation that if there exists a color $u$ that is distinct from $y$, $a$, or $x$ that is between $y$ and $a$ in $S$, then the sandwich $yuy$ blocks $x$ in $S^{\prime}R$, a contradiction. ∎
The following lemma is a crucial piece of our inductive proof, because if (I) and (IV) do not hold for a sock ordering $X$ which satisfies the hypothesis, then, because (III) implies that $X$ is sortable, we can determine if a sock ordering is sortable and exactly how to sort $X$ by considering a proper subordering.
@@
Let $X$ be a sock ordering and let $a$ and $d$ be distinct colors in $X$. Let $\overline{\theta}$ be a nonempty sorting sequence for $X$ and let $(S,R):=\overline{\theta}(\emptyset,X)$. Let $d$ be the last color in $\overline{\theta}$. Furthermore, assume there exists a substack $S^{\prime}\subseteq S$ and subordering $R^{\prime}\subseteq R$ such that the color $a$ is beneath the top color on $S^{\prime}$, $a$ appears in $R^{\prime}$, and $a$ is the only good sortable color in $S^{\prime}R^{\prime}$. Then, one of the following is true:
@@
    $X-d$ is sortable and if a sequence $\theta$ fully sorts $X-d$, then when we attempt to sort $X$ with the sequence $\theta$, there is a stage where the color $d$ only appears as the top color on the stack (this condition implies that $X$ is sortable); or
@@
Assume that (I), (II), and (III) do not hold. We will show that (IV) must hold. Let $x$ be the top color on $S^{\prime}$ and $D\subseteq X$ be such that $X=YDW$ where the first element of $D$ is the first $d$ in $X$ and the final element of $D$ is the final $d$ in $X$. Note that the subordering $R$ must follow the last $d$ in $X$, so $R\subseteq W$. Similarly, everything in $S$ must precede the last $d$ in $X$, so $S\subseteq YD$. Since $R\subseteq W$ and $a\in R^{\prime}\subseteq R$, we have $a\in W$, so the fact there are no $d$-sandwiches that block $a$ in $X$ implies that the only colors that could appear in $D$ are $a$ and $d$. In particular, the color $x$ does not appear in $D$, so the fact that $a$ is beneath $x$ on $S^{\prime}$ and $S^{\prime}\subseteq S\subseteq YD$ implies that $ax\subseteq Y$. Note that because $d$ is sorted by $\overline{\theta}$ but $a$ and $x$ are not, there is no sandwich that consists only of $a$ and $x$ that blocks $d$ in $X$. Therefore, the color $a$ does not appear in $D$ as otherwise the fact that $ax\subseteq Y$ would imply the existence of a sandwich of the form $axa$ before the final $d$ in $X$. As such, $d$ is the only color that appears in $D$ and we can assume without loss of generality that the color $d$ only appears once in $X$, so $D=d$.
@@
By assumption, the statements (II) and (III) are both false, so there exists a sorting sequence $\theta$ that fully sorts $X-d=YW$ such that when we apply $\theta$ to $X=YDW=YdW$ the color $d$ never appears only as the top color on the stack. Let $\theta_{a}$ be the list of initial colors in $\theta$ up to, but not including, $a$ (so $\theta_{a}$ is empty if and only if $\theta$ begins with $a$). Note that no color in $S^{\prime}R^{\prime}$ could appear before $a$ in $\theta$ because $a$ is the only good sortable color in $S^{\prime}R^{\prime}\subseteq YW$ and $\theta$ fully sorts $YW$. We claim that this implies that no color that appears after $ax$ in $YW$ is in $\theta_{a}$. To see this, note that sorting such a color would put $ax$ onto the stack, but that would make it impossible to sort $a$ before sorting $x$. In particular, because $ax\subseteq Y$, this means that no color that appears in $W$ is in $\theta_{a}$. Therefore, $a$ is the color in $W$ that is sorted first by $\theta$, and $\theta_{a}$ is a sorting sequence for $X=YdW$. Furthermore, if we let $(S_{1},R^{\prime}_{1})=\theta_{a}(YW)$, then $(S_{1},R_{1})=\theta_{a}(YdW)$ where $R^{\prime}_{1}=R_{1}-d$. Therefore, because $a$ is sortable in $(S_{1},R^{\prime}_{1})$ and no $d$-sandwich blocks $a$ in $X$, the color $a$ is also sortable in $(S_{1},R_{1})$. Recall that no color that appears after $ax$ is in $\theta_{a}$ and $x$ is not in $\theta_{a}$. This, with the fact that $axda\subseteq SdR\subseteq X$, implies that $xda\subseteq R_{1}$. Therefore, sorting $a$ in the state $(S_{1},R_{1})$ pushes $xd$ onto the stack. Let $y$ be the color that immediately precedes $d$ in $R_{1}$ (so we could have $y=x$). Note that since $ax\subseteq Y$ and either $x=y$ or $xy\subseteq Y$, in all cases, we have
$ay\subseteq Y.$ (1)
@@
Recall that, by the previous arguments, $d$ is pushed on the stack when we attempt to apply $\theta$ to $X$, so $\theta$ cannot be a sorting sequence for $X$ as otherwise we would have $(d,\emptyset)=\theta(X)$, a contradiction to the fact that the color $d$ never appears as the top color on the stack. Therefore, there must exist a color in $\theta$ that causes the stack to be unsorted when $\theta$ is applied to $X$. Let $z$ be the first such color and let $\theta_{z}$ be the list of initial colors in $\theta$ up to but not including $z$. That is, $\theta_{z}$ is the longest initial sequence of $\theta$ that is a sorting sequence for $X$. Let $(S_{2},R_{2}):=\theta_{z}(X)$. Note that by our previous arguments, $a$ must appear in $\theta_{z}$ and $a$ is the color in $W$ that is sorted first by $\theta_{z}$ and $d\notin W$. By Lemma 4, we have
$W\supseteq I(a,t_{0},\dotsc,t_{n})R_{2}=t_{0}aI(t_{0},\dotsc,t_{n})R_{2},$ (2)
where $t_{0},\dotsc,t_{n}$ is the sequence of colors that appear above $d$ on the stack after $a$ is sorted.